A box contains 3 white and 8 yellow balls. If two balls are drawn in succession without replacement, what is the probability that: a.) both are white? b.) the first ball is white and the second ball is yellow?
Answer by ewatrrr(24785)


Hi
Thanks for the 'complete' scoop on this one. 
11 balls: 3W , 8Y
two balls drawn in succession 'without replacement'

P(2 White) =  (3/11)(2/10) = .0545

P(1st W & 2nd Y) = (3/11)(8/10) = .2182

Wish You the Best in your Studies.